OpenStudy (anonymous):

show your work: 3/(x^2) + 2/x = 10/ (2x^2)

OpenStudy (ash2326):

We have \[\frac 3{x^2}+\frac 2 x=\frac{10}{2x^2}\] Multiply whole equation by 2x^2 cause it's the LCM of x, x^2 and 2x^2 so we get \[2x^2\times \frac 3{x^2}+2x^2\times \frac 2 x=2x^2\times \frac{10}{2x^2}\] we get now \[6+2x\times 2=10\] or we get \[6+4x=10\] Can you solve for x now?

OpenStudy (ash2326):

@Elysse2012 do you get this?

OpenStudy (anonymous):

how did you get from multiplying both sides by 2x^2 to the equation 6+2x x 2 = 10?

OpenStudy (anonymous):

oh wait... nvm. thank you!

OpenStudy (ash2326):

got it?

OpenStudy (anonymous):

yup